# Fedor V. Fomin

### v1

@@ 2.1 Spanning Lemma
We first deal with highly-connected graphs. More precisely, we demonstrate that, given an injective mapping $f:V(H)\to V(G)$, we can either compute a TM-embedding of $H$ in $G$ respecting $f$ and spanning $V(G)$, or find an independent set of size $k$ in $G$; this holds if the connectivity of $G$ is sufficiently high. This step follows the approach of Jedlicková and Kratochvíl [JK23, Theorem 7], who proved that for every $k>1$ and $\ell>1$, if a graph $G$ satisfies $\alpha(G)<k$ and $G$ is $\max\{k\ell,10\ell\}$-connected, then $G$ is $-\ell$-linked. Recall that a graph $G$ is $\ell$-linked if it has at least $2\ell$ vertices and for every set of $\ell$ disjoint pairs $(s_{j},t_{j})$ of distinct vertices, there exist $\ell$ vertex-disjoint paths in $G$ such that the endpoints of the $j$-path are $s_{j}$ and $t_{j}$, for each $j\in[\ell]$. A graph is called $\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
To adapt the approach of Jedlicková and Kratochvíl for our purposes, we need to address two issues. The first issue is minor; we need to slightly strengthen the connectivity requirement to simplify the case analysis and allow for possible coinciding endpoints of the paths (which is not allowed in [JK23]). The second issue is more serious—the proof of Jedlicková and Kratochvíl is not constructive, meaning that it cannot be used to construct the linkage. However, we need a procedure constructing the linkage for computing a TM-embedding of $H$. The proof of Jedlicková and Kratochvíl builds on the following result of Thomas and Wollan [TW05], which is also non-constructive.
@@
Proof of Claim 1. Since $G$ is $2\ell$-connected, by Menger’s theorem [Men27, Die17], there are $2\ell$ internally vertex-disjoint paths connecting vertices of $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ with distinct vertices of $C$. These paths can be constructed in the standard fashion via a network flow algorithm; in particular, using the Ford-Fulkerson algorithm [FF56] to find $\ell$-flow in $\mathcal{O}(\ell\cdot|G|)$ running time. Then for each $i\in[\ell]$, we have two paths, one connecting $s_{i}$ with $c_{2i-1}$ and the other connecting $t_{i}$ with $c_{2i}$ for some $c_{2i-1},c_{2i}\in C$. Note that it is possible that one or two of these paths are trivial, i.e., $s_{i}=c_{2i-1}$ or $t_{i}=c_{2i}$ can be true. As $C$ is a clique in $G$ and $c_{2i-1}c_{2i}\in E(G)$, these paths can be concatenated, resulting in the $s_{i}t_{i}$-path $P_{i}$. Since $c_{1},c_{2},\ldots,c_{2\ell}$ are pairwise distinct, any two paths among $P_{1},P_{2},\ldots,P_{\ell}$ do not share any common vertices. We obtain $\ell$ paths as required. ∎
@@
Proof of Claim 3. Note that $G$ necessary has either an independent set of size $r$ or a clique of size $s$ because the corresponding Ramsey number satisfies $R(r,s)\leq\binom{r+s-2}{r-1}$ [ES09]. The algorithm is recursive. Since $r,s>0$, $|V(G)|>1$ and $G$ has at least one vertex, denote it by $v$. If $r=1$ or $s=1$, then algorithm outputs $\{v\}$ and stops. For $r,s\geq 2$, consider vertex sets $A=N_{G}(v)$ and $B=V(G)\setminus\{v\}\setminus A$. Since $|V(G)|=|A|+|B|+1$ and $V(G)\geq\binom{r+s-2}{r-1}$ we have that either $|A|\geq\binom{r+s-3}{r-1}$ or $|B|\geq\binom{r+s-3}{r-2}$. If $|A|\geq\binom{r+s-3}{r-1}$ then $G[A]$ has either an independent set of size $r$ or a clique of size $s-1$. In this case, the algorithm makes a recursive call to itself for the input $G[A]$, $r>0$ and $s-1>0$. If its output is a clique of size $s-1$, the algorithm additionally appends $v$ to it, resulting in a clique of size $s$. Thus, the algorithm obtains either an independent set of size $r$, or a clique of size $s$ as required. The case of $|B|\geq\binom{r+s-3}{r-2}$ is symmetric, as $v$ can be appended to any independent set in $G[B]$. Because the values of $\binom{x+y-2}{y-1}=\binom{x+y-2}{x-1}$ for $x\leq r$ and $y\leq s$ can be computed in $\mathcal{O}(rs)$ time using the recursive formula for binomial coefficients, we obtain that the running time of the algorithm is $\mathcal{O}(n^{2})$ where $n=|V(G)|$. The proof of the claim is complete. ∎
@@
  • If $|V(G)|\leq(k+2\ell)^{k}$, the algorithm applies the subroutine of 2 to find the solution, outputs it and stops.
  • Otherwise, since $(k+2\ell)^{k}\geq\binom{k+2\ell-2}{k-1}$, the algorithm applies the subroutine of 3 with $r=k$ and $s=2\ell$.
@@
  • Otherwise, the output is a clique of size $2\ell$ in $G$. The algorithm uses the subroutine of 1 to find the solution, then outputs it and stops.
@@
First, the algorithm reduces to the case where $H$ has no isolated vertices. Since each isolated vertex $v$ in $H$ has a fixed image $f(v)$ in $G$, the algorithm simply removes $v$ from $H$ and $f(v)$ from $G$. In the following steps of the algorithm, if a spanning TM-embedding of $H$ is produced, the algorithm extends this embedding with the removed images of isolated vertices. Such an embedding remains spanning for $G$, while all described procedures are performed in polynomial time. Note that removing an equal number of vertices from both $G$ and $H$ preserves the connectivity constraint in the statement. From now on, every vertex of $H$ has at least one incident edge, while $G$ is ($\max\{k+2,10\}\cdot|E(H)|$)-connected. In order to proceed, it is convenient to reformulate the problem in terms of disjoint paths: Given $\ell$ vertex pairs $(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{\ell},t_{\ell})$ of distinct vertices in $G$, find a family of internally-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$ such that $P_{i}$ connects $s_{i}$ with $t_{i}$ for each $i\in[\ell]$ and each vertex of $G$ belongs to at least one path, and the union of the paths spans $V(G)$. To see that the reformulation is equivalent, put $\ell=|E(H)|$ and $(s_{i},t_{i})=(f(x_{i}),f(y_{i}))$, where $x_{i},y_{i}$ are the endpoints of the $i$-th edge of $H$. Then, given the paths $P_{i}$, it is straightforward to construct the subgraph $M$ by taking the union of the paths; $(M,f)$ then is the desired TM-embedding of $H$ in $G$. The reverse is also possible: given a TM-embedding $(M,f)$ of $H$ in $G$ spanning $V(G)$, the respective path can be found by traversing $M$ from $f(x_{i})$ to $f(y_{i})$. From now on, we do not refer to TM-embeddings, but work with the equivalent disjoint-path statement above. The first step of the algorithm is to find the initial family of paths $P_{1},P_{2},\ldots P_{\ell}$, that satisfies all constraints except for spanning $V(G)$. That is, $P_{i}$ is an $s_{i}t_{i}$-path, and for each $i\neq j$, $P_{i}$ and $P_{j}$ have no common vertices except for possibly their endpoints. For this, the algorithm invokes the algorithm of Lemma 1 as a subroutine. The call is valid since $G$ is $10\ell$-connected, therefore in time $2^{(k+\ell)^{\mathcal{O}(k)}}+|G|^{\mathcal{O}(1)}$ the initial sequence $P_{1},P_{2},\ldots,P_{\ell}$ is produced. The only step left is to ensure that the spanning constraint is satisfied by the path sequence. The algorithm achieves this incrementally. That is, the algorithm enlarges one of the paths in the sequence without violating any other constraints. If we show how this can be done exhaustively and in polynomial time, it is clearly sufficient to find the desired TM-embedding.
@@
If $|S|\leq(k+2)\ell$, then two of these paths end in $s_{1}$ and $v$, where $v$ is the neighbor of $s_{1}$ in $P_{1}$. Denote these paths by $Q_{1}$ and $Q_{2}$. These two paths are internally disjoint, both starting in $x$ and ending in $s_{1}$ and $v$, respectively. The algorithm enlarges $P_{1}$ by replacing it with the path $(s_{1}Q_{1}x)\circ(xQ_{2}v)\circ(vP_{1}t_{1})$, that is, $x$ becomes embedded in $P_{1}$ between $s_{1}$ and $v$. Here and next, for a path $P$ and vertices $u$, $v$ on $P$ we denote by $uPv$ the subpath of $P$ from $u$ to $v$, and for two paths $P_{1}$ and $P_{2}$ sharing an endpoint, we denote by $P_{1}\circ P_{2}$ their concatenation. This operation strictly increases the span of the path sequence. It is left for the algorithm to process the case $(k+2)\ell<|S|$. There are precisely $(k+2)\ell$ paths that were found by the Ford-Fulkerson algorithm, so at least $k\ell$ paths end in internal vertices of $P_{1},P_{2},\ldots,P_{\ell}$. Hence, there is $i\in[\ell]$ such that at least $k$ paths end in internal vertices of $P_{i}$. The algorithm takes arbitrary $k$ of them. Denote the $k$ selected paths by $Q_{1},Q_{2},\ldots,Q_{k}$ and their respective endpoints in $P_{i}$ by $v_{1},v_{2},\ldots,v_{k}$. For each $a\in[k]$, let $u_{a}$ be the successor of $v_{a}$ in $P_{i}$ if one follows vertices on $P_{i}$ from $s_{i}$ to $t_{i}$. If the set $\{u_{1},u_{2},\ldots,u_{k}\}$ is an independent set in $G$, then the algorithm reports this independent set and stops. Otherwise, there is an edge $u_{a}u_{b}$ in $G$ for some $a,b\in[k]$. Without loss of generality, $v_{a}$ goes before $v_{b}$ on $P_{i}$ in the order from $s_{i}$ to $t_{i}$. The algorithm then enlarges $P_{i}$ by replacing it with $(s_{i}P_{i}v_{a})\circ(v_{a}Q_{a}x)\circ(xQ_{b}v_{b})\circ(v_{b}P_{i}u_{a})\circ u_{a}u_{b}\circ(u_{b}P_{i}s_{i})$. Again, $x$ is embedded into $P_{i}$ while no original vertices of $P_{i}$ are lost, increasing the total span. The algorithm repeats the described procedure until paths $P_{1},P_{2},\ldots,P_{\ell}$ span the whole vertex set of $G$, or an independent set of size $k$ is encountered. This procedure is repeated at most $|V(G)|$ times, so the total running time of this part is polynomial in the size of the instance. The correctness of the algorithm follows from the discussion. The proof of the lemma is complete. ∎
@@ 2.2 Merging Lemma
The main idea of the proof is to enumerate all possible interactions of the desired TM-model with $S$. We encapsulate the choice of the interaction in an object we call a cut descriptor, to be formally defined next. Intuitively, a cut descriptor captures the behavior of a TM-model of $H$ in $G$ in the subset $S$ and its neighborhood: Given a TM-model $(M,T)$ of $H$ in $G$, its respective cut descriptor is obtained from $M$ by dissolving all non-terminals outside of $S$, where vertices not in $S$ are represented implicitly by the index of their component. The latter part is crucial to keep the number of choices the function of the parameter. As we will see later, such an implicit representation is sufficient since Lemma 2 always provides a spanning TM-model (or an independent set) and allows to fix an arbitrary injection of the terminals.
Formally, a cut descriptor is the tuple $(W,M_{W},\xi,f_{W})$, where $W$ is a set of vertices (indexed implicitly from $1$ to $|W|$), $M_{W}$ is a graph with the vertex set $S^{\prime}\cup W$ with $S^{\prime}\subseteq S$, $\xi:W\to[t]$ is a mapping from $W$ to indices of the components $C_{1}$, …, $C_{t}$, and $f_{W}:V(H)\to S^{\prime}\cup W$ is an injective mapping from $V(H)$ to vertices of $M_{W}$. Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $W_{i}=\{w\in W:\xi(w)\in C_{i}\}$. We require the following properties for $(W,M_{W},\xi,f_{W})$ to be a cut descriptor:
@@ 5. CD5
Let $(M,f)$ be a list TM-embedding of $H$ in $G$. Then there exist a cut descriptor $(W,M_{W},\xi,f_{W})$ such that $M_{W}[S]=M[S]$, and for each $i\in[t]$, $M$ hits $C_{i}$ if and only if $M_{W}$ hits $W_{i}$, touches $C_{i}$ if and only if $M_{W}$ touches $W_{i}$ and $|V(M)\cap C_{i}|=|W_{i}|$, and avoids $C_{i}$ if and only if $W_{i}=\emptyset$.
Proof of Claim 5. We construct the desired cut descriptor as follows. First, let $W$ be the subset of vertices in $G$ that are either terminals in $V(G)\setminus S$ or neighbors of $S$ in the model $M$, i.e., $W=(\operatorname{Im}f\setminus S)\cup N_{M}(S)$. We then dissolve all vertices in $V(M)\setminus(S\cup W)$, which are all non-terminals by definition of $W$, to obtain $M_{W}$ from $M$. We set $f_{W}$ as the respective mapping of the terminals obtained from $f$. Clearly, CD1 is fulfilled for $(M_{W},T_{W})$, and CD2 also holds immediately by construction. We set $\xi$ so that $W_{i}=W\cap C_{i}$ for each $i\in[t]$. We now proceed to verify the remaining properties of a cut descriptor.
For CD3, consider $w,w^{\prime}\in W$ with $\xi(w)\neq\xi(w^{\prime})$. Let $w\in W_{i},w^{\prime}\in W_{j}$ in $M_{W}$, $i\neq j$, and $w\in C_{i}$, $w^{\prime}\in C_{j}$ in $G$. Since $S$ separates $C_{i}$ and $C_{j}$ in $G$, and $M$ is a subgraph of $G$, $w$ and $w^{\prime}$ are not adjacent in $M$. By construction of $M_{W}$, only non-terminals outside of $S$ and $W$ are dissolved; every such non-terminal lies in $C_{i^{\prime}}$ together with its neighbors, for some $i^{\prime}\in[t]$, and dissolving it only changes edges between vertices of $C_{i^{\prime}}$. Therefore, $w$ and $w^{\prime}$ are also not adjacent in $M_{W}$.
@@
For the last claim of the lemma, we argue again by construction: since the dissolving operation only affects edges inside $C_{i}$, for some $i$, and dissolving a vertex in $C_{i}$ leaves $C_{i}$ with at least one edge inside, the status of the component (hit/touched/avoided) is always preserved between $M$ and $M_{W}$. Specifically, if an edge of $M$ was present in $C_{i}$, then after the dissolving, an edge is still present in $M_{W}$ in $W_{i}$. If no edge but some vertices of $M$ are present in $C_{i}$ then all these vertices are terminals or neighbors of $S$, and $|M\cap C_{i}|=|W_{i}|$, and if no vertices of $M$ are in $C_{i}$, then also $W_{i}$ is empty. ∎
Now, the algorithm proceeds as follows. First, it branches over the choice of a cut descriptor. Then, for a fixed cut descriptor, the algorithm picks an arbitrary assignment fulfilling CD5. Since the assignment fixes the vertices of $W$ in $G$, we run the algorithm of Lemma 2 on each $C_{i}$ to find a TM-model connecting the vertices of $W_{i}$ as prescribed by $M_{W}$. If any of the runs returns an independent set of size $k$, we output it and stop. Otherwise, we augment the model $(M_{W},T_{W})$ to a model $(M,T)$ of $H$ in $G$ via the returned model in each $C_{i}$. Out of all models of $H$ in $G$ obtained in the branches, the algorithm returns the one of maximum size. The above algorithm is shown in detail in Algorithm 1.
@@
1 $\mathcal{M}\leftarrow\emptyset$; for $|W|\leftarrow 0$ to $h+2s$ do
2       $W\leftarrow$vertex set of size $|W|$; foreach $M_{W}$ a graph on $S\cup W$, $\xi:W\to[t]$, $f_{W}:V(H)\to S\cup W$ do
3             verify CD1CD4, verify CD5 by a matching instance and pick arbitrary suitable $\tau$; if any of CD1CD5 fail then
@@
6            $M\leftarrow M_{W}$, $f\leftarrow f_{W}$; for $i\leftarrow 1$ to $t$ do
7                   $T_{i}\leftarrow\{w\in W:\xi(w)=i\}$, $H_{i}=M_{W}[T_{i}]$; if $H_{i}$ contains no edges then

### v2

@@ 4.1 Spanning Lemma
We first deal with highly-connected graphs. More precisely, we demonstrate that, given an injective mapping $f:V(H)\to V(G)$, we can either compute a TM-embedding of $H$ in $G$ respecting $f$ and spanning $V(G)$, or find an independent set of size $k$ in $G$; this holds if the connectivity of $G$ is sufficiently high. Recall that a graph $G$ is $\ell$-linked if it has at least $2\ell$ vertices and for every set of $\ell$ disjoint pairs $(s_{j},t_{j})$ of distinct vertices, there exist $\ell$ vertex-disjoint paths in $G$ such that the endpoints of the $j$-path are $s_{j}$ and $t_{j}$, for each $j\in[\ell]$. A graph is called $-\ell$-linked if, in addition, the set of $\ell$ paths covers all vertices of $G$.
We need a procedure constructing the linkage for computing a TM-embedding of $H$. For that we need a constructive version of the following theorem of Thomas and Wollan [TW05], which proof is non-constructive.
@@
Proof of Claim 8. Since $G$ is $2\ell$-connected, by Menger’s theorem [Men27, Die17], there are $2\ell$ internally vertex-disjoint paths connecting vertices of $T=\bigcup_{i=1}^{\ell}\{s_{i},t_{i}\}$ with distinct vertices of $C$. These paths can be constructed in the standard fashion via a network flow algorithm; in particular, using the Ford-Fulkerson algorithm [FF56] to find $\ell$-flow in $\mathcal{O}(\ell\cdot|G|)$ running time. Then for each $i\in[\ell]$, we have two paths, one connecting $s_{i}$ with $c_{2i-1}$ and the other connecting $t_{i}$ with $c_{2i}$ for some $c_{2i-1},c_{2i}\in C$. Note that it is possible that one or two of these paths are trivial, i.e., $s_{i}=c_{2i-1}$ or $t_{i}=c_{2i}$ can be true.
As $C$ is a clique in $G$ and $c_{2i-1}c_{2i}\in E(G)$, these paths can be concatenated, resulting in the $s_{i}t_{i}$-path $P_{i}$. Since $c_{1},c_{2},\ldots,c_{2\ell}$ are pairwise distinct, any two paths among $P_{1},P_{2},\ldots,P_{\ell}$ do not share any common vertices. We obtain $\ell$ paths as required. ∎
@@
Proof of Claim 10. Note that $G$ necessary has either an independent set of size $r$ or a clique of size $s$ because the corresponding Ramsey number satisfies $R(r,s)\leq\binom{r+s-2}{r-1}$ [ES09]. The algorithm is recursive. Since $r,s>0$, $|V(G)|>1$ and $G$ has at least one vertex, denote it by $v$. If $r=1$ or $s=1$, then algorithm outputs $\{v\}$ and stops. For $r,s\geq 2$, consider vertex sets $A=N_{G}(v)$ and $B=V(G)\setminus\{v\}\setminus A$. Since $|V(G)|=|A|+|B|+1$ and $V(G)\geq\binom{r+s-2}{r-1}$ we have that either $|A|\geq\binom{r+s-3}{r-1}$ or $|B|\geq\binom{r+s-3}{r-2}$. If $|A|\geq\binom{r+s-3}{r-1}$ then $G[A]$ has either an independent set of size $r$ or a clique of size $s-1$. In this case, the algorithm makes a recursive call to itself for the input $G[A]$, $r>0$ and $s-1>0$. If its output is a clique of size $s-1$, the algorithm additionally appends $v$ to it, resulting in a clique of size $s$. Thus, the algorithm obtains either an independent set of size $r$, or a clique of size $s$ as required. The case of $|B|\geq\binom{r+s-3}{r-2}$ is symmetric, as $v$ can be appended to any independent set in $G[B]$.
Because the values of $\binom{x+y-2}{y-1}=\binom{x+y-2}{x-1}$ for $x\leq r$ and $y\leq s$ can be computed in $\mathcal{O}(rs)$ time using the recursive formula for binomial coefficients, we obtain that the running time of the algorithm is $\mathcal{O}(n^{2})$ where $n=|V(G)|$. The proof of the claim is complete. ∎
@@
  • If $|V(G)|\leq(k+2\ell)^{k}$, the algorithm applies the subroutine of 9 to find the solution, outputs it and stops.
  • Otherwise, since $(k+2\ell)^{k}\geq\binom{k+2\ell-2}{k-1}$, the algorithm applies the subroutine of 10 with $r=k$ and $s=2\ell$.
@@
  • Otherwise, the output is a clique of size $2\ell$ in $G$. The algorithm uses the subroutine of 8 to find the solution, then outputs it and stops.
@@
First, the algorithm reduces to the case where $H$ has no isolated vertices. Since each isolated vertex $v$ in $H$ has a fixed image $f(v)$ in $G$, the algorithm simply removes $v$ from $H$ and $f(v)$ from $G$. In the following steps of the algorithm, if a spanning TM-embedding of $H$ is produced, the algorithm extends this embedding with the removed images of isolated vertices. Such an embedding remains spanning for $G$, while all described procedures are performed in polynomial time.
Note that removing an equal number of vertices from both $G$ and $H$ preserves the connectivity constraint in the statement. From now on, every vertex of $H$ has at least one incident edge, while $G$ is ($\max\{k+2,10\}\cdot|E(H)|$)-connected.
In order to proceed, it is convenient to reformulate the problem in terms of disjoint paths: Given $\ell$ vertex pairs $(s_{1},t_{1}),(s_{2},t_{2}),\ldots,(s_{\ell},t_{\ell})$ of distinct vertices in $G$, find a family of internally-disjoint paths $P_{1},P_{2},\ldots,P_{\ell}$ such that $P_{i}$ connects $s_{i}$ with $t_{i}$ for each $i\in[\ell]$ and each vertex of $G$ belongs to at least one path, and the union of the paths spans $V(G)$.
To see that the reformulation is equivalent, put $\ell=|E(H)|$ and $(s_{i},t_{i})=(f(x_{i}),f(y_{i}))$, where $x_{i},y_{i}$ are the endpoints of the $i$-th edge of $H$. Then, given the paths $P_{i}$, it is straightforward to construct the subgraph $M$ by taking the union of the paths; $(M,f)$ then is the desired TM-embedding of $H$ in $G$. The reverse is also possible: given a TM-embedding $(M,f)$ of $H$ in $G$ spanning $V(G)$, the respective path can be found by traversing $M$ from $f(x_{i})$ to $f(y_{i})$. From now on, we do not refer to TM-embeddings, but work with the equivalent disjoint-path statement above.
The first step of the algorithm is to find the initial family of paths $P_{1},P_{2},\ldots P_{\ell}$, that satisfies all constraints except for spanning $V(G)$. That is, $P_{i}$ is an $s_{i}t_{i}$-path, and for each $i\neq j$, $P_{i}$ and $P_{j}$ have no common vertices except for possibly their endpoints. For this, the algorithm invokes the algorithm of Lemma 3 as a subroutine. The call is valid since $G$ is $10\ell$-connected, therefore in time $2^{(k+\ell)^{\mathcal{O}(k)}}+|G|^{\mathcal{O}(1)}$ the initial sequence $P_{1},P_{2},\ldots,P_{\ell}$ is produced.
The only step left is to ensure that the spanning constraint is satisfied by the path sequence. The algorithm achieves this incrementally. That is, the algorithm enlarges one of the paths in the sequence without violating any other constraints. If we show how this can be done exhaustively and in polynomial time, it is clearly sufficient to find the desired TM-embedding.
@@
If $|S|\leq(k+2)\ell$, then two of these paths end in $s_{1}$ and $v$, where $v$ is the neighbor of $s_{1}$ in $P_{1}$. Denote these paths by $Q_{1}$ and $Q_{2}$. These two paths are internally disjoint, both starting in $x$ and ending in $s_{1}$ and $v$, respectively. The algorithm enlarges $P_{1}$ by replacing it with the path $(s_{1}Q_{1}x)\circ(xQ_{2}v)\circ(vP_{1}t_{1})$, that is, $x$ becomes embedded in $P_{1}$ between $s_{1}$ and $v$. Here and next, for a path $P$ and vertices $u$, $v$ on $P$ we denote by $uPv$ the subpath of $P$ from $u$ to $v$, and for two paths $P_{1}$ and $P_{2}$ sharing an endpoint, we denote by $P_{1}\circ P_{2}$ their concatenation. This operation strictly increases the span of the path sequence.
It is left for the algorithm to process the case $(k+2)\ell<|S|$. There are precisely $(k+2)\ell$ paths that were found by the Ford-Fulkerson algorithm, so at least $k\ell$ paths end in internal vertices of $P_{1},P_{2},\ldots,P_{\ell}$. Hence, there is $i\in[\ell]$ such that at least $k$ paths end in internal vertices of $P_{i}$. The algorithm takes arbitrary $k$ of them. Denote the $k$ selected paths by $Q_{1},Q_{2},\ldots,Q_{k}$ and their respective endpoints in $P_{i}$ by $v_{1},v_{2},\ldots,v_{k}$. For each $a\in[k]$, let $u_{a}$ be the successor of $v_{a}$ in $P_{i}$ if one follows vertices on $P_{i}$ from $s_{i}$ to $t_{i}$.
If the set $\{u_{1},u_{2},\ldots,u_{k}\}$ is an independent set in $G$, then the algorithm reports this independent set and stops. Otherwise, there is an edge $u_{a}u_{b}$ in $G$ for some $a,b\in[k]$. Without loss of generality, $v_{a}$ goes before $v_{b}$ on $P_{i}$ in the order from $s_{i}$ to $t_{i}$. The algorithm then enlarges $P_{i}$ by replacing it with $(s_{i}P_{i}v_{a})\circ(v_{a}Q_{a}x)\circ(xQ_{b}v_{b})\circ(v_{b}P_{i}u_{a})\circ u_{a}u_{b}\circ(u_{b}P_{i}t_{i})$. Again, $x$ is embedded into $P_{i}$ while no original vertices of $P_{i}$ are lost, increasing the total span.
The algorithm repeats the described procedure until paths $P_{1},P_{2},\ldots,P_{\ell}$ span the whole vertex set of $G$, or an independent set of size $k$ is encountered. This procedure is repeated at most $|V(G)|$ times, so the total running time of this part is polynomial in the size of the instance.
The correctness of the algorithm follows from the discussion. The proof of the lemma is complete. ∎
@@ 4.2 Merging Lemma
Figure 3: An illustration to the proof of Lemma 5: A graph $G$ with a TM-model of $H$, the graph $M_{W}$ of the cut descriptor corresponding to the model, and the graph $H$, the model of which is sought. The images of the vertices of $H$ in $G$ and $M_{W}$ that are fixed by $f$ and $f_{W}$, respectively, are indicated by the matching node shapes.
The main idea of the proof is to enumerate all possible interactions of the desired TM-model with $S$. We encapsulate the choice of the interaction in an object we call a cut descriptor, to be formally defined next. Intuitively, a cut descriptor captures the behavior of a TM-model of $H$ in $G$ in the subset $S$ and its neighborhood: Given a TM-model $(M,T)$ of $H$ in $G$, its respective cut descriptor is obtained from $M$ by dissolving all non-terminals outside of $S$, where vertices not in $S$ are represented implicitly by the index of their component. The latter part is crucial to keep the number of choices the function of the parameter. As we will see later, such an implicit representation is sufficient since Lemma 4 always provides a spanning TM-model (or an independent set) and allows to fix an arbitrary injection of the terminals. See Figure3 for an illustration of the relationship between a TM-model of $H$ in $G$ and its cut descriptor.
Formally, a cut descriptor is the tuple $(W,M_{W},\xi,f_{W})$, where $W$ is a set of vertices (indexed implicitly from $1$ to $|W|$), $M_{W}$ is a graph with the vertex set $S^{\prime}\cup W$ with $S^{\prime}\subseteq S$, $\xi:W\to[t]$ is a mapping from $W$ to indices of the components $C_{1}$, …, $C_{t}$, and $f_{W}:V(H)\to S^{\prime}\cup W$ is an injective mapping from $V(H)$ to vertices of $M_{W}$. Denote $T_{W}=\operatorname{Im}f_{W}$, and for each $i\in[t]$, denote $T_{i}=\{w\in W:\xi(w)\in C_{i}\}$. We require the following properties for $(W,M_{W},\xi,f_{W})$ to be a cut descriptor:
@@ 5. CD5
Let $(M,f)$ be a list TM-embedding of $H$ in $G$. Then there exist a cut descriptor $(W,M_{W},\xi,f_{W})$ such that $M_{W}[S]=M[S]$, and for each $i\in[t]$, $M$ hits $C_{i}$ if and only if $M_{W}$ hits $T_{i}$, touches $C_{i}$ if and only if $M_{W}$ touches $T_{i}$ and $|V(M)\cap C_{i}|=|T_{i}|$, and avoids $C_{i}$ if and only if $T_{i}=\emptyset$.
Proof of Claim 12. We construct the desired cut descriptor as follows. First, let $W$ be the subset of vertices in $G$ that are either terminals in $V(G)\setminus S$ or neighbors of $S$ in the model $M$, i.e., $W=(\operatorname{Im}f\setminus S)\cup N_{M}(S)$. We then dissolve all vertices in $V(M)\setminus(S\cup W)$, which are all non-terminals by definition of $W$, to obtain $M_{W}$ from $M$. We set $f_{W}$ as the respective mapping of the terminals obtained from $f$. Clearly, CD1 is fulfilled for $(M_{W},T_{W})$, and CD2 also holds immediately by construction. We set $\xi$ so that $T_{i}=W\cap C_{i}$ for each $i\in[t]$. We now proceed to verify the remaining properties of a cut descriptor.
For CD3, consider $w,w^{\prime}\in W$ with $\xi(w)\neq\xi(w^{\prime})$. Let $w\in T_{i},w^{\prime}\in W_{j}$ in $M_{W}$, $i\neq j$, and $w\in C_{i}$, $w^{\prime}\in C_{j}$ in $G$. Since $S$ separates $C_{i}$ and $C_{j}$ in $G$, and $M$ is a subgraph of $G$, $w$ and $w^{\prime}$ are not adjacent in $M$. By construction of $M_{W}$, only non-terminals outside of $S$ and $W$ are dissolved; every such non-terminal lies in $C_{i^{\prime}}$ together with its neighbors, for some $i^{\prime}\in[t]$, and dissolving it only changes edges between vertices of $C_{i^{\prime}}$. Therefore, $w$ and $w^{\prime}$ are also not adjacent in $M_{W}$.
@@
For the last claim of the lemma, we argue again by construction: since the dissolving operation only affects edges inside $C_{i}$, for some $i$, and dissolving a vertex in $C_{i}$ leaves $C_{i}$ with at least one edge inside, the status of the component (hit/touched/avoided) is always preserved between $M$ and $M_{W}$. Specifically, if an edge of $M$ was present in $C_{i}$, then after the dissolving, an edge is still present in $M_{W}$ in $T_{i}$. If no edge but some vertices of $M$ are present in $C_{i}$ then all these vertices are terminals or neighbors of $S$, and $|M\cap C_{i}|=|T_{i}|$, and if no vertices of $M$ are in $C_{i}$, then also $T_{i}$ is empty. ∎
Now, the algorithm proceeds as follows. First, it branches over the choice of a cut descriptor. Then, for a fixed cut descriptor, the algorithm picks an arbitrary assignment fulfilling CD5. Since the assignment fixes the vertices of $W$ in $G$, we run the algorithm of Lemma 4 on each $C_{i}$ to find a TM-model connecting the vertices of $T_{i}$ as prescribed by $M_{W}$. If any of the runs returns an independent set of size $k$, we output it and stop. Otherwise, we augment the model $(M_{W},T_{W})$ to a model $(M,T)$ of $H$ in $G$ via the returned model in each $C_{i}$. Out of all models of $H$ in $G$ obtained in the branches, the algorithm returns the one of maximum size. The above algorithm is shown in detail in Algorithm 1.
@@
1
$\mathcal{M}\leftarrow\emptyset$;
3for $|W|\leftarrow 0$ to $h+2s$ do
4       $W\leftarrow$vertex set of size $|W|$;
5      foreach $M_{W}$ a graph on $S\cup W$, $\xi:W\to[t]$, $f_{W}:V(H)\to S\cup W$ do
6             verify CD1CD4, verify CD5 by a matching instance and pick arbitrary suitable $\tau$;
7            if any of CD1CD5 fail then
@@
10
11            $M\leftarrow M_{W}$, $f\leftarrow f_{W}$;
12            for $i\leftarrow 1$ to $t$ do
13                   $T_{i}\leftarrow\{w\in W:\xi(w)=i\}$, $H_{i}=M_{W}[T_{i}]$;
14                  if $H_{i}$ contains no edges then